two numbers are in the ratio 8:11 and their sum is 57. find the numbers
Answer by Alan3354(69443)

two numbers are in the ratio 8:11 and their sum is 57. find the numbers
8 + 11 = 19
57/19 = 3
---------
3*8 = 24
3*11 = 33
---------
8:1::24:33